OpenStudy (anonymous):

Use quadratic formula to solve for y 3x^2- 2(sqrt3)xy +y^2 +6(sqrt3)x -2y -40=0

myininaya (myininaya):

so we can put y in terms of x?

myininaya (myininaya):

\[(1)y^2+(-2 \sqrt{3}x-2)y+(3x^2+6 \sqrt{3}x-40)=0\]

myininaya (myininaya):

\[\text{ where }a=1, b=-2 \sqrt{3}-2 , c=3x^2+6 \sqrt{3} x-40\]

myininaya (myininaya):

\[y=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

well the option that is closest to what i get when i plug those in is y=1+sqrt3(x) +/- sqrt(-4sqrt3(x) =41)

myininaya (myininaya):

\[y=\frac{-(-2 \sqrt{3}-2) \pm \sqrt{(-2 \sqrt{3}-2)^2-4(1)(3x^2+6 \sqrt{3}x-40)}}{2(1)}\]

myininaya (myininaya):

this is crazy lookin' lol

myininaya (myininaya):

\[y=\frac{2 \sqrt{3} +2 \pm \sqrt{4(3)-2(-2 \sqrt{3} (2))+2^2-12x^2-24 \sqrt{3} x+160}}{2}\]

OpenStudy (anonymous):

haha yes i know! that is what i got when i plugged them in. However, i have multiple choice, and t hat is not an option. \[y= 1+ \sqrt{3}x \pm \sqrt{-4\sqrt{3}x +41}\]

OpenStudy (anonymous):

this is the only similar option

myininaya (myininaya):

\[y=\frac{2 \sqrt{3} +2 \pm \sqrt{12+8 \sqrt{3}+4-12x^2-24 \sqrt{3} x+160}}{2}\]

OpenStudy (anonymous):

idk its not one of my options :(

myininaya (myininaya):

\[y=\frac{2 \sqrt{3}+2 \pm \sqrt{-12x^2-24 \sqrt{3}x+176+8 \sqrt{3}}}{2}\] i'm still simplying i'm just showing my steps

myininaya (myininaya):

simplifying*

OpenStudy (anonymous):

oh okay! sorry!

myininaya (myininaya):

\[y=\frac{2 \sqrt{3}+2 \pm \sqrt{4} \sqrt{-3x^2-6 \sqrt{x} +44+2 \sqrt{3}}}{2}\]

myininaya (myininaya):

\[y=\frac{2 \sqrt{3} +2 \pm 2 \sqrt{-3x^2-6 \sqrt{x}+44+2 \sqrt{3}}}{2}\]

myininaya (myininaya):

notice each term has the common factor 2

myininaya (myininaya):

\[y=\frac{ \sqrt{3}+1 \pm 1 \sqrt{-3x^2-6 \sqrt{x}+44+2 \sqrt{3}}}{1}\] \[y= \sqrt{3}+1 \pm \sqrt{-3x^2-6 \sqrt{x}+44+2 \sqrt{3}}\]

myininaya (myininaya):

oops i'm missing x in my b term

myininaya (myininaya):

:(

myininaya (myininaya):

i have to do this on paper lol

myininaya (myininaya):

can you give me a sec?

OpenStudy (anonymous):

its alright! of course i can! do you think when you are done you could help me with another similar problem?

myininaya (myininaya):

ok i got it

myininaya (myininaya):

i'm gonna scan it

OpenStudy (anonymous):

ok:)

myininaya (myininaya):

which is what you thought gj

OpenStudy (anonymous):

yes! thank you so much

myininaya (myininaya):

that is weird the file posted twice

myininaya (myininaya):

how many times do you see the file?

myininaya (myininaya):

it keeps posting :(

OpenStudy (anonymous):

like 5 at least haha

myininaya (myininaya):

ok i have to do something post your new question on the left ok and if i get back i will look at it

OpenStudy (anonymous):

okay! i will.

myininaya (myininaya):

:) fun problem by the way

OpenStudy (anonymous):

haha yes it was! thanks for your help!